OpenStudy (anonymous):

Evaluate without using a calculator: A.) sin(-390)-tan 495 B.) sin90 cos135 - tan30

OpenStudy (cwrw238):

A sin (-390) = sin(-390 + 360) = sin(-30) = - sin 30 tan 495 = tan (495 - 360) = tan 135 = -tan45 use the special triangles to find these values

OpenStudy (triciaal):

A. hint -390 + 360 = - 30 and 495 = 360 + 90 + 45

OpenStudy (cwrw238):